Connect Problem CP 9-7 (algo) Fill in the blanks in the table below. The problem is a puzzle so the blanks are not necessarily filled in sequentially Instructions: Round your answers to the nearest whole number. Enter your answers without a comma (thousands separator) Average Product Marginal Product Labor Total Output 10 20 30 40 50 60 200 800 1,800 3,200 5,000 7,200 20 40 60 80 100 120 20 60 100 180
Answer #1

Average products

For Q = 20, AP = TP/Q = 800/20 = 40 units

For Q = 60, AP = TP/Q = 7200/60 = 120 units

Total products

For Q = 30, TP = AP x Q = 60 x 30 = 1800 units

For Q = 50, TP = AP x Q = 100 x 50 = 5000 units

Marginal products

For Q = 40, MP = (3200 - 1800)/10 = 140 units

For Q = 60, MP = (7200 - 5000)/10 = 220 units